# A Heavy Seed Black Hole Mass Function at High Redshift - Prospects for LISA

### v1

@@ 3.4 Growth Through Mergers
Having analysed the SNR from the instantaneous merger case we now move onto determining the merger timescales for a more realistic case i.e. where the MBHs take some additional time to merge following the host halo merger. Using the equations described in §2.7 we compute the delay times for the MBH mergers, for MBHs seeded with both scaling relations and both growth methods, and display them in Figure 7. As before the Rarepeak region is displayed in the top panel with the Normal region in the lower panel, for all the growth models we calculated in this analysis. On the x-axis we show the mass of the primary black hole and on the y-axis the merger time. Each merger candidate is coloured by the initial separation.
What we see from Figure 7 is that almost half of the mergers have delay times exceeding the Hubble time particularly for the lower mass end of the MBH mass spectrum. There is also a very clear function of separation in the determining the merger timescales. The initial separation between the black holes is calculated by taking the minimum of the distance between the black holes in the snapshot before the host halos are considered merged and the radius of the merged halo. For the higher MBH masses which are involved in a merger (i.e. those with masses in excess of at least $10^{3}\ \rm{M}_{\odot}$) we see that dynamical friction can drag these MBH to coalescence in less than a Hubble time and with delay times at low as 100 Myr. Our calculations also show a number of light seeds with masses less than $10^{3}\ \rm{M}_{\odot}$ with delay times less than a Hubble time. These lighter seed mergers have short delay times because of the small separations that exist between them just after the galactic merger.
@@ 3.6 Comparing our Number Density Rates to the Existing Literature
Habouzit et al. (2016) similarly found that number densities of order $10^{-6}$ cMpc-3 at z $\sim 10$ result from a threshold Jcrit = 300 when analysing output from several cosmological hydrodynamical simulations that cover a large range of box sizes and resolutions. Lessening the metallicity constraint to incorporate metal-poor rather than metal-free haloes improves the number densities somewhat (Dunn et al., 2018; Chon & Omukai, 2020; Regan et al., 2020) bringing number density values up to values of approximately $10^{-4}$ cMpc-3 at z $\sim 10$ (Visbal, Haiman & Bryan, 2014b; Lupi, Haiman & Volonteri, 2021; Trinca et al., 2022; Spinoso et al., 2023).
@@
The values we find are at least two to three orders of magnitude higher than the more optimistic LW channel results (and in fact are closer to some semi-analytic models based on very efficient growth of light seeds (e.g. Spinoso et al., 2023)). Our values are founded on the rapid growth of the host halo predominantly following on from the results of Wise et al. (2019). A comparable study by Lupi, Haiman & Volonteri (2021) used a semi-analytical treatment on top of n-body only calculations of a quasar progenitor environment to calculate number densities in a similar way. Lupi, Haiman & Volonteri (2021) found number density values of approximately 0.1 cMpc-3 at $z\sim 10$ (after accounting for the impact of their overdense realisation). This value compares extremely well to our estimate (see again Figure 9).
@@ 3.7 LISA Event Rates from Mergers of Heavy Seed Massive Black Holes
3.8 Observation of Renaissance-like black holes using Electromagnetic and Gravitational Wave signatures
@@
While LISA will be able to detect high redshift mergers, the constraint on the redshift and mass parameters of these mergers will prove difficult. There is little known about the dark matter distribution of the Universe at high redshifts, therefore the effect that this distribution will have on the gravitational wave signatures detected is unknown and therefore large uncertainties will have to be taken into account. These uncertainties can be slightly alleviated if we can manage to preform multimessenger observations of the mergers. For example the X-ray telescope Athena (Barcons et al., 2012), which is due to operate at the same time as LISA, will aim to use its higher resolution to preform accurate measurements of accretion signatures onto black holes. For example Mangiagli et al. (2024) argue that combining measurements from LISA and Athena will allow us to accurately study mergers up to a redshift of $z=8$ with the brightest sirens.
The challenge of detecting this high redshift black hole population will prove substantial in the upcoming years. The combination of the numerous upcoming gravitational wave and electromagnetic observation missions should provide a more comprehensive analysis of the number density of mid to high mass black holes in the early Universe.
@@ 4 Summary & Conclusions
In this paper, we set out to study the halo environments potentially conducive to heavy seeds formation in the early Universe via post process modelling in the Renaissance simulations. In order for a halo to host a heavy seed, it would need to satisfy four criteria: (i) reaching a mass needed for atomic cooling, (ii) a high inflow rate ($\dot{M}>0.1$ $\rm{M_{\odot}\ yr^{-1}}$ ), (iii) low metallicity ($Z<10^{-3}\rm{\ Z_{\odot}}$) and (iv) compactness. In effect we find that the high mass inflow rate is the strictest criteria with the other criteria more easily attained.
@@
Finally, MBH mergers are also included in our analysis both from the point of view of growth but also from the point of view of observability with future gravitational wave observatories. For analysing MBH mergers we again consider two scenarios. In the naive case we assume that the BBHs merge instantly as part of the host halo merger and secondly we consider the case where the black holes must coalesse via dynamical friction. The latter case can, in some instances, result in significant delay times between the host galaxy merger and the BBH merger. Our main conclusions are:
@@
    We calculate LISA event rates of both the Normal and Rarepeak regions and find that the high redshift regime will provide numerous candidates for LISA detection in its proposed 10 year lifespan, with $\sim 80$ mergers expected to be observed using the merger predictions from the Normal region (with $\sim 10^{3}$ mergers predicted from the RP region). If LISA is able to detect the signals from these mergers, we will be able to further constrain the formation processes of the black hole populations at this epoch. Since the uncertainties on the parameter constraints on GW signals alone will be significantly high, combined EM measurements using telescopes like Athena will be needed to further constrain the redshifts and redshifted masses of merger events in the early Universe.

### v2

@@ 3.4 Growth Through Mergers
Having analysed the SNR from the instantaneous merger case we now move onto determining the merger timescales for a more realistic case i.e. where the MBHs take some additional time to merge following the host halo merger. Using the equations described in §2.7 we compute the delay times for the MBH mergers, for MBHs seeded with both scaling relations and both growth methods, and display them in Figure 10. As before the Rarepeak region is displayed in the top panel with the Normal region in the lower panel, for all the growth models we calculated in this analysis. On the x-axis we show the mass of the secondary black hole and on the y-axis the merger time. Each merger candidate is coloured by the initial separation.
What we see from Figure 10 is that almost half of the mergers have delay times exceeding the Hubble time particularly for the lower mass end of the MBH mass spectrum. There is also a very clear function of separation in the determining the merger timescales. The initial separation between the black holes is calculated by taking the minimum of the distance between the black holes in the snapshot before the host halos are considered merged and the radius of the merged halo. For the higher MBH masses which are involved in a merger (i.e. those with masses in excess of at least $10^{3}\ \rm{M}_{\odot}$) we see that dynamical friction can drag these MBH to coalescence in less than a Hubble time and with delay times at low as 100 Myr. Our calculations also show a number of light seeds with masses less than $10^{3}\ \rm{M}_{\odot}$ with delay times less than a Hubble time. These lighter seed mergers have short delay times because of the small separations that exist between them just after the galactic merger. It must be noted that the delay time here is based on the parameters and structure of the halo at the time of merger. As the halo evolves through cosmic time it will grow and change in structure, thus affecting the delay time for the mergers.
@@
We also note that as the host halos grow through mergers with smaller halos, there is an increase in the Star Formation Rate in the halo, as shown in Chen et al. (2014). Provided that the halo is massive enough to overcome the effects of stellar feedback, the amount of gas accretion onto the halo also increases. The replenishing of this gas reservoir provides an opportunity for the massive black hole that a halo could host, to gain a boost in the accretion from this gas.
@@ 3.6 Comparing our Number Density Rates to the Existing Literature
Habouzit et al. (2016) similarly found that number densities of order $10^{-6}$ cMpc-3 at z $\sim 10$ result from a threshold Jcrit = 300 when analysing output from several cosmological hydrodynamical simulations that cover a large range of box sizes and resolutions. Lessening the metallicity constraint to incorporate metal-poor rather than metal-free haloes improves the number densities somewhat (Dunn et al., 2018; Chon & Omukai, 2020; Regan et al., 2020) bringing number density values up to approximately $10^{-4}$ cMpc-3 at z $\sim 10$ (Visbal, Haiman & Bryan, 2014b; Lupi, Haiman & Volonteri, 2021; Trinca et al., 2022; Spinoso et al., 2023).
@@
The values we find, through appealing to the so-called ‘rapid assembly model’ (e.g. Yoshida et al., 2003; Wise et al., 2019), are at least two to three orders of magnitude higher than the more optimistic LW channel results (and in fact are closer to some semi-analytic models based on very efficient growth of light seeds (e.g. Trinca et al., 2022; Spinoso et al., 2023)). As discussed in §XYZ our values are founded on the rapid growth of the host halo predominantly following on from the results of Wise et al. (2019). A comparable study by Lupi, Haiman & Volonteri (2021) used a semi-analytical treatment on top of n-body only calculations of a quasar progenitor environment to calculate number densities in a similar way. Lupi, Haiman & Volonteri (2021) found number density values of approximately 0.1 cMpc-3 at $z\sim 10$ (after accounting for the impact of their overdense realisation). This value compares extremely well to our estimate (see again Figure 12).
@@ 3.7 LISA Event Rates from Mergers of Heavy Seed Massive Black Holes
3.8 Detection of High-z Black Holes and Comparison with Other Semi-analytical Models
@@
Other semi-analytical and simulation models have explored the extent to which GW observatories will explore the the high-z population of massive black holes (e.g. Sesana, Volonteri & Haardt, 2007; Tanaka & Haiman, 2009; Dayal et al., 2019; Valiante et al., 2020). The general consensus is that if we are able to constrain the parameters from black hole mergers and analyse a merger rate (whose predicted values range from $10^{1}-10^{3}\rm{yr}^{-1}$ with LISA) using the synergy of various gravitational wave detectors, we should be able to constrain the massive black hole formation models which dominate the early Universe.
@@ 4 Summary & Conclusions
In this paper, we set out to study the halo environments potentially conducive to heavy seeds formation in the early Universe via post process modelling in the Renaissance simulations. In order for a halo to host a heavy seed, it would need to satisfy four criteria: (i) reaching a mass needed for atomic cooling, (ii) a high inflow rate ($\dot{M}>0.1$ $\rm{M_{\odot}\ yr^{-1}}$ ), (iii) low metallicity ($Z<10^{-3}\rm{\ Z_{\odot}}$) and (iv) compactness. In effect we find that the high mass inflow rate is the strictest criteria with the other criteria more easily attained. Because the final redshifts of each region are quite high, they don’t fully capture the end of massive black hole formation, as can be seen in fig.3, therefore we are underestimating the mergers that will happen after the end times of Renaissance. Nonetheless it is important to include these mergers in our black hole formation models as they will contribute to the signal that LISA will detect.
@@
Finally, MBH mergers are also included in our analysis both from the point of view of growth but also from the point of view of observability with future gravitational wave observatories. For analysing MBH mergers we again consider two scenarios. In the naive case we assume that the BBHs merge instantly as part of the host halo merger and secondly we consider the case where the black holes must coalesce via dynamical friction. The latter case can, in some instances, result in significant delay times between the host galaxy merger and the BBH merger. Our main conclusions are:
@@
    We calculate LISA event rates of both the Normal and Rarepeak regions and find that the high redshift regime will provide numerous candidates for LISA detection in its proposed 10 year lifespan, with $\sim 80$ mergers expected to be observed using the merger predictions from the Normal region (with $\sim 10^{3}$ mergers predicted from the RP region). If LISA is able to detect the signals from these mergers, we will be able to further constrain the formation processes of the black hole populations at this epoch.